OpenStudy (beth12345):

A frictionless disc of mass 0.50kg is moving in a straight line across an air table at a speed of 2.4m/s when it bumps into an elastic band stretched between two fixed posts. If the elastic band exerts an average opposing force of 1.2N on the disc for 1.5s, what will be the final velocity of the disc?

OpenStudy (beth12345):

The answer I was given is... -1.2m/s, but I don't know how to get it.

OpenStudy (anonymous):

force x time =final momentum - initial momentum, find the initial momentum of the disc. you have the force and time, exereted by the band.

OpenStudy (beth12345):

ok, just a sec

OpenStudy (beth12345):

is it 1.2?

OpenStudy (anonymous):

didn't you get the answer?

OpenStudy (beth12345):

I don't know what the final momentum is

OpenStudy (anonymous):

when you think of a physics question, just go think about how it is happening. The disc e with some mass (m) is moving with some velocity, (v1) right. It keeps going like that until it reaches the elastic band and it causes it to change its velocity to (v2). That means that the momentum changes as well. from mv1 to mv2. Now, we can calculate the change in momentum,(force x time, look at the equation above), that is exerted by the band. so the initial momentum is mv1 the final momentum is mv2 the change in momentum is Force x time does this help?

OpenStudy (beth12345):

oh ok, so am I finding the change in momentum by using (Force x time)?

OpenStudy (anonymous):

force x time =(final momentum - initial momentum) --->change in momentum \[force \times time= \Delta p\]

OpenStudy (beth12345):

k, I got\[\Delta p = 1.8\]

OpenStudy (anonymous):

this means that a force 1.2 N was applied on the disc by the band for the time of 1.5 s, that caused it to change momentum, (change the velocity)

OpenStudy (anonymous):

yup, that's right

OpenStudy (beth12345):

would I get the -1.2 value by 2.4-1.8=0.6 and then 1.8-0.6=1.2?

OpenStudy (anonymous):

I think you should focus more on trying to understand the question than finding the answer. and no, force x time = final momentum - initial momentum force x time= mv2 -mv1

OpenStudy (beth12345):

I don't understand what to do next

OpenStudy (anonymous):

force x time= mv2 -mv1 put in the values force=1.2N time=1.5s m=0.5kg v1=2.4m/s v2= find that!

OpenStudy (beth12345):

thanks :)